# Dynamical correlation effects in twisted bilayer graphene under strain and lattice relaxation

### v1

@@ Abstract
Figure 3: a) Sketch of the bare (zero-hybridization) dispersion of the $f$- and $c$-electrons in the relaxed THF model (left) and the corresponding density of states (right), b) Inverse charge compressibility $\partial\mu/\partial\nu$ as a function fo filling in TBLG with and without lattice relaxation and $0.15\%$ strain compared with experimental data [93, 102, 91]. P-symmetry breaking due to lattice relaxation results in stronger peaks and troughs in the inverse compressibility on the electron-doped side.
.4 Entropy with $\nu$
@@
Entropy has been recently used to experimentally assess the presence and degeneracy of local moments in TBLG  [70]. We calculate the entropy as a function of filling, for different values of temperature, by means of the Maxwell relation
@@
Fig. 4 shows the entropy of TBG with and without strain and relaxation at two temperatures. At $11$ K, the entropy curves with and without strain are qualitatively different. In the unstrained case, the entropy curve follows a rough semi-circular envelope spanning $\nu\in[-4,4]$ with a maximum at the charge neutrality point. In contrast, in the strained case, the entropy goes to zero at the charge neutrality point, with two semi-circular envelopes spanning $\nu\in[-4,0]$ and $\nu\in[0,4]$. This can be explained by the splitting of the $8$-fold degenerate flat-band manifold into two $4$-fold degenerate anti-bonding and bonding bands by strain. The vanishing entropy at the charge neutrality point is a consequence of the gap between the anti-bonding and bonding flat bands. At higher temperatures ($30$ K), thermal excitation can overcome the strain gap restoring the $8$-fold degeneracy of the local moments. The degeneracy is reflected in the fine features of the entropy curve. There are local maxima between neighboring integer fillings giving rise to an overall $8$-lobed structure. Consequently, at high temperatures, both unstrained and strained data have the same qualitative features: $8$ lobes and finite entropy at CNP. The shape of the entropy curves match favorably with experimental measurements [70, 90].
@@
Figure 4: The entropy as a function of filling of unperturbed TBLG (orange lines) compared to that of strained ($0.15\%$) and relaxed TBLG (blue lines). At high temperature compared to $\Delta_{\mathrm{strain}}$ (dashed lines), the entropy is 88-peaked (44-peaked for $\nu\in[0,4]$) for both strained and unstrained cases. At low temperature, only the $f_{-}$ electrons are active, leading to a decrease in entropy at small doping in the strained case. This difference is absent in the unstrained case. The errorbars (shaded area) are obtained by propagating the relative mean square deviation of a sample of DMFT-CTQMC iterations after convergence. In the unstrained case, the entropy on the hole-doping side is determined using $P$-symmetry and reflecting about the CNP (thin lines).
.5 Conclusions
We have performed a comprehensive set of DMFT simulations for the THF model in the presence of strain and lattice relaxation. We have shown through spectral and statistical indicators that the inclusion of strain and lattice relaxation terms into the DMFT treatment reproduces key experimental features that had been missing in previous efforts, giving rise to a categorical improvement in the predictive power of the model. We identify the splitting of the flat-band manifold by strain to be the key driver behind two phenomena identified in recent experiments—the presence of a persistent feature in the local spectral density corresponding to excitations at circa +10+10 and 10-10 meV on the hole- and electron-doped side respectively [89], and the transition from an $8$-fold degenerate local moment to a $4$-fold degenerate local moment upon lowering the temperature [70]. We find that at low temperatures, one of the two strain-split sectors is inactive, and energetically placed so as to precisely account for the persistent feature in QTM and STM experiments. The freezing of the inactive sector gives a direct explanation for the lowered degeneracy of the local moment, and our numerical estimate of the entropy gives an accurate match with recent experimental data. Our observation of the freezing of the inactive charge sector has also implications for the future modeling of correlations in TBLG, as projecting onto the strain-split basis could allow techniques such as numerical renormalization group to be used in a controlled approximation [101, 80]. We also incorporate P-symmetry breaking terms into our calculations by including the lattice relaxation terms, which results in the bonding flat bands being more dispersive than the anti-bonding flat bands. We find that this asymmetry carries over to the interacting properties, with stronger variations of the charge compressibility on the electron-doped side than on the hole-doped side.
.6 Acknowledgements
@@
: G.R., L.C., and T.W. acknowledge support from the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ – EXC 2056 (Project No. 390715994), and SPP 2244 (WE 5342/5-1 project No. 422707584). L.C. gratefully acknowledges the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b158cb. G.R., L.C., R.V., G.S., and T.W. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through QUAST FOR 5249 (Project No. 449872909, projects P4 and P5). R.V. thanks the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the TRR 288 - 422213477 and project Nr. VA 117/23-1 — 509751747. B.A.B. and H. H. were supported by the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070), the Office of Naval Research (ONR Grant No. N00014-20-1-2303), the Global Collaborative Network Grant at Princeton University, the Simons Investigator Grant No. 404513, the NSF-MERSEC (Grant No. MERSEC DMR 2011750), Simons Collaboration on New Frontiers in Superconductivity (SFI-MPS- NFS-00006741-01), the Schmidt Foundation at Princeton University and the Princeton Catalyst Initiative. H. H. and D. C. also acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833).
@@
Supplemental material for Dynamical correlation effects in twisted bilayer graphene under strain and lattice relaxation
@@
The noninteracting Hamiltonian of the THF model, as originally derived in [31], has the form
@@
To this Hamiltonian we add the following terms, a subset of those derived in [30], which account for uniaxial heterostrain
@@
The THF Hamiltonian involves a sum over moiré lattice sites and and momenta alike. We consider a single lattice site labeled by $\mathbf{R}=(0,0)$, and restrict our $k$-summation on the first mBZ, sampled with a regular centered grid as in [78]. The $k$-points on the edges and corners of the hexagonal BZ are degenerate, and are sampled with a relative weight of $1/2$ and $1/3$ respectively. We then strain the obtained $k$-point mesh with the transformation
@@
These are density-density interactions acting within the $f$ subspace, between $f$ and $c$ and within $c$ respectively. The remaining terms, such as the fcf-c exchange interaction, have been neglected since they do not qualitatively alter the resulting picture in the symmetric phase [78]. We define the density operators
@@
We assume for simplicity the same interaction coefficient for the $\Gamma_{3}$ and $\Gamma_{1}\oplus\Gamma_{2}$ dispersive orbitals, so that $W_{a}=W=47.12meV$, and $V=48.33meV$. These assumptions are consistent with the study of the unstrained and unrelaxed model in the symmetric phase in [78, 79], and do not alter the qualitative picture.
@@
We now show that a similar conclusion can be alternatively reached using symmetry arguments in combination with a perturbative expansion in the strength of the strain. The most general strain term in the THF model contains the isotropic strain parameter $\epsilon_{+}$, the anisotropic $\epsilon_{-}$ and the shear $\epsilon_{xy}$. These transform as $x^{2}+y^{2}$, $x^{2}-y^{2}$ and $xy$ respectively under spatial symmetry. $f$ electron forms the $\Gamma_{3}$ irreducible representation [31, 30]. We expand the local hybridization function in powers of the strength of strain, and discuss the symmetry properties of each term. Without loss of generality, we focus on valley $\eta=+$ and spin $s$. The local hybridization can be written as
@@
For the numerical solution of the THF model we employ the w2dynamics software suite [104], which belongs to the class of DMFT-CTQMC solvers. While extremely robust and able to work with continuous baths and in an extended temperature range, this class of solvers famously suffers from a limitation called fermionic sign problem (see, e.g. Ref. [105]). No exact prescription for the improvement of the quality of the Monte Carlo sign exists, but it is generally true that off-diagonal local impurity Hamiltonians and hybridization functions tend to aggravate the problem [106]. In addition to a generally better Monte Carlo sign, another advantage of working with orbital/valley/spin-diagonal impurity problems is the necessity of evaluating a smaller number of matrix elements for the local Green’s function and self-energy. Furthermore, upon postprocessing there is no need to analytically continue the off-diagonal elements of $G$ and $\Sigma$, which is a particularly delicate process.
@@
In the rotated basis, the $f$-subspace consists of doubly-degenerate levels, whose local splitting is $2M_{f}\epsilon_{-}$. This splitting is constant in temperature and linear in strain, as described in [30] and visualized in Fig. S2.
@@
We solve the strained and relaxed THF model within the framework of Dynamical Mean-Field theory [107], making use of the Continuous-Time Quantum Monte Carlo (CTQMC) solver w2dynamics [104]. As discussed in section S1.3, the intraction Hamiltonian is treated at two different levels of approximation: the $H_{U}$ is dynamically accounted for within the DMFT approximation, while the remaining ones are mean-field decoupled. Concordantly, the electronic self-energy of the lattice problem is obtained self-consistently: first, starting from the density matrix at each loop of the DMFT simulation, the mean-field terms are obtained. These are summed to the non-interacting Hamiltonian. The $f$-orbital peojection of this operator represents the impurity Hamiltonian which is solved via CTQMC. The resulting self-energy is upfolded in the combined $f\oplus c$ space by padding it with zeros, since no dynamical correlation effects are considered for the dispersive bands. Self-consistent update of the density matrix and self-energy are then performed until convergence. Since we treat all interaction terms apart from $H_{U}$ at the Hartree level, we can perform a further simplification [79] that allows us to describe them as an effective “double-counting” term, i.e. an orbital-selective chemical potential acting on the correlated subspace only:
@@
We run the simulations, consistently with [78], on 144 CPU cores, using ${\texttt{Nmeas}}=7.5\cdot 10^{4}$ measurements per core for each Monte Carlo simulation, with a number of steps ${\texttt{Ncorr}}=2000$ per core between successive measurements to avoid autocorrelation effects and ${\texttt{Nwarmup}}=2\cdot 10^{6}$ warmup steps per core before each simulation. We measured in imaginary time domain and used the Legendre polynomial basis, with maximum order $\texttt{NLegMax}=40$.
@@
Analytical continuation is performed starting from the last DMFT run using the MaxEnt method from the ana_cont software suite [108]. Once the self-energy and Green’s function are analytically continued, they can be rotated back in the original orbital basis, where $H^{ff}$ is not diagonal. We used a regular mesh of $1001$ real-axis frequency points in the interval [250:+250][-250:+250] meV for $A(\omega)$ and [300:+300][-300:+300] meV for $\Sigma(\omega)$ and no preblur.
@@
Fig.  S6 shows the momentum-resolved spectral function along the $K_{M}\rightarrow\Gamma_{M}\rightarrow-K_{M}\rightarrow M_{M}$ high-symmetry path. The panels in the first row refer to all integer fillings in the range $\nu\in[-4;+4]$ The doping values have been chosen to be directly comparable to the experimental data in [89]. The simulations are in a rather good agreement with the QTM results: they feature flat bands almost everywhere in the mBZ, apart from near the $\Gamma_{M}$ point where they become dispersive. At the CNP, the flat bands are separated by about 35 meV, in agreement with the experiment, and they touch near the $\Gamma_{M}$ point. At positive and negative integer $\nu$ values, a persistence spectral maximum due to the flat band is visible at negative and positive $\omega$ respectively, the value of which is in the range $\pm[10;20]$ meV. Increasing occupation from $\nu=-4$, the lower flat band first remains rather pinned at zero frequency up to $\nu=-1$, after which it shifts to lower frequency and becomes the negative-bias persistent feature for electron doping. Comparing positive and negative doping of the same size, it is interesting to notice how the electron-doped side features a larger spectral gap at large doping with respect to the hole-doped side: this is most evident for $\nu=+4$, where the system is in a clear band-insulating state. This is to be compared to the $\nu=-4$ case, which features a vanishing indirect gap at the Fermi level, and is a product of the $P$-symmetry breaking relaxation terms, which shift the $f$-bands asymmetrically within the hybridization gap. The $\nu=-4$ case is particularly delicate, since the presence or absence of the gap is directly linked to the model parameters, in particular the effective chemical potentials deriving from lattice relaxation.
Figure S7: Occupation with respect to CNP as a function of the chemical potential, total value (a) and orbital-valley resolved (b-c), for T=11.6 K. The blue curves refer to DMFT simulations in presence of strain and lattice relaxation, while the orange one are the unperturbed results from [78].
@@
One of the innovations provided by the Quantum Twisting Microscope is the possibility of accessing the spectral function at different points in the Brillouin Zone. By probing $A(\omega)$ at the $\Gamma_{M}$ point, the authors of [89] observe a spectral maximum inside the flat band gap, which evolves with doping in a manner reminiscent of the $\mu(\nu)$ curve (see Fig.1e) in the main text.
@@
In Fig. S9, in the left panel we show the spectral function at the $\Gamma_{M}$ point as a function of doping. This features four wiggling spectral maxima, that evolve from positive to negative frequency as a function of doping. The other panels of Fig. S9 show the momentum-resolved spectral functions for $c$-electrons at five different values of doping, including the CNP. The positions and relative distances of the peaks, are again represented by colored circles and dashed lines. It is immediate to notice that the circles all refer to local spectral maxima of $c$ character. By comparing the relative distances, it is also easy to notice that the four peaks are almost rigidly shifted upon varying the chemical potential. This fact can be understood in the following way: since the $c$ electrons are essentially uncorrelated, their behavior of their spectral distribution is well approximated by that of noninteracting electronic bands. Varying the chemical potential by $\Delta\mu$ has the only effect of rigidly shifting the whole band structure by $\Delta\mu$ with respect to zero frequency. This will follow the peculiar wiggling behavior that results from the $f/c$ spectral weight reshuffling, sometimes referred to as Dirac revivals, as described in [78].

### v2

@@ Abstract
Figure 3: a Sketch of the bare (zero-hybridization) dispersion of the $f$- and $c$-electrons in the relaxed THF model (left) and the corresponding density of states (right), b Inverse charge compressibility $\partial\mu/\partial\nu$ as a function fo filling in TBLG with and without lattice relaxation and $0.15\%$ strain and $T=11.6$ K, compared with experimental data [pierceUnconventionalSequenceCorrelated2021, zondinerCascadePhaseTransitions2020, saitoIsospinPomeranchukEffect2021]. P-symmetry breaking due to lattice relaxation results in stronger peaks and troughs in the inverse compressibility on the electron-doped side.
At very low T, symmetry-broken integer-filling insulators are more stable on the electron side [stepanovUntyingInsulatingSuperconducting2020, saitoIndependentSuperconductorsCorrelated2020]. Lattice relaxation effects are responsible for particle-hole symmetry breaking, as Hartree-Fock studies have shown [kwanKekuleSpiralOrder2021, wagnerGlobalPhaseDiagram2022]. Yet, mean-field solutions are unable to correctly capture the sawtooth behavior of compressibility and the negative compressibility regions [raiDynamicalCorrelationsOrder2024, saitoIsospinPomeranchukEffect2021, zondinerCascadePhaseTransitions2020, pierceUnconventionalSequenceCorrelated2021].
To explicitly quantify the impact of lattice relaxation induced $P$-symmetry breaking on theormodynamic observables, we study the relation between the chemical potential and the total filling. Fig. 2(b) shows the inverse charge compressibility, $\partial\mu/\partial\nu$, of the strained and relaxed THF model compared to the unperturbed THF model from  [raiDynamicalCorrelationsOrder2024] and experimental data [zondinerCascadePhaseTransitions2020, pierceUnconventionalSequenceCorrelated2021, saitoIsospinPomeranchukEffect2021].
We find that the addition of strain and lattice relaxation terms allows us to capture an important qualitative feature missing in the unperturbed THF model. Multiple experiments [zondinerCascadePhaseTransitions2020, pierceUnconventionalSequenceCorrelated2021, saitoIsospinPomeranchukEffect2021] across different samples consistently find a strong particle-hole asymmetry: the variations in the charge compressibility are markedly stronger on the e-doped side compared to the h-doped side. The THF model with strain and lattice relaxation reproduces this phenomenon, and it can be understood by considering the primary effect of the lattice relaxation part of the Hamiltonian, which is a relative downward shift of the $f$-electron energies with respect to the charge neutrality point of the $c$ Dirac bands. This is shown schematically in Fig. 3(a) where the bare $f$ dispersion is represented by the red line, shifted down with respect to the touching point of the $c$ bands. As a consequence, there is a higher $c$ density of states available on the hole-doped side compared to the electron-doped side, and therefore there is stronger hybridization on the hole-doped side compared to the electron-doped side. This suppresses the local moment on the hole-doped side relative to the electron-doped side, which results in the softening of the compressibility variations. This is consistent with Hartree-Fock simulations which find that the ordered insulators on the electron-doped side have a larger gap than the ones on the hole-doped side [kwanKekuleSpiralOrder2021, wagnerGlobalPhaseDiagram2022]. Note that one might naively expect that since the upper flat band is more dispersive, correlations would be stronger on the hole-doped side. However, the crucial factor affecting the strength of correlations is the hybridization, which in turn correlates positively with the higher $c$ density of states available on the hole-doped side.
Another key improvement with respect to previous simulations regards the position and shape of the inverse compressibility maxima. Though different experimental results show variations in the position and relative size of the peaks for electron doping (positive $\nu$), the behavior for hole doping (negative $\nu$) is remarkably consistent [saitoIsospinPomeranchukEffect2021, rozenEntropicEvidencePomeranchuk2021, pierceUnconventionalSequenceCorrelated2021, zondinerCascadePhaseTransitions2020], and shows a marked depinning of the inverse compressibility maxima from integer total occupations. The inclusion of strain and relaxation terms immediately leads to a better estimation of the position and size of the compressibility maxima on the hole-doping side (blue solid line) with respect to the unstrained model (dashed black line in Fig. 3b). The increased hybridization on the hole-doped side again plays a crucial role in the depinning. The hybridization has a smoothening effect making the $f$-sector less quantum-dot like, lowering the penalty of non-integer occupation. In other words, the system is more metallic on the hole-doped side than on the electron-doped side (see further discussion in the supplemental material [refsupplement]).
.5 Entropy
@@
Entropy has been recently used to experimentally assess the presence and degeneracy of local moments in TBLG  [zhangHeavyFermionsMass2025]. We calculate the entropy as a function of filling, for different values of temperature, by means of the Maxwell relation
@@
Figure 4: Entropy as a function of filling of unperturbed TBLG (orange lines) compared to that of strained ($0.15\%$) and relaxed TBLG (blue lines). The error bars (shaded area) are obtained by propagating the relative mean square deviation of a sample of DMFT-CTQMC iterations after convergence. In the unstrained case, the entropy on the hole-doping side is not calculated directly, since it is is determined by $P$-symmetry. It is denoted by thin lines without error bars.
Fig. 4 shows the entropy of TBG with and without strain and relaxation at two temperatures. At $11$ K, the entropy curves with and without strain are qualitatively different. In the unstrained case, the entropy curve oscillates around a rough semi-circular envelope spanning $\nu\in[-4,4]$ with a maximum at the charge neutrality point. In contrast, in the strained case, the entropy goes to zero at the charge neutrality point, with two semi-circular envelopes spanning $\nu\in[-4,0]$ and $\nu\in[0,4]$. This can be explained by the splitting of the $8$-fold degenerate flat-band manifold into two $4$-fold degenerate anti-bonding and bonding bands by strain. The vanishing entropy at the charge neutrality point is a consequence of the gap between the anti-bonding and bonding flat bands. At higher temperatures ($30$ K), thermal excitation can overcome the strain gap restoring the $8$-fold degeneracy of the local moments. The degeneracy is reflected in the fine features of the entropy curve. There are local maxima between neighboring integer fillings giving rise to an overall $8$-lobed structure. Consequently, at high temperatures, both unstrained and strained data have the same qualitative features: $8$ lobes and finite entropy at CNP. The shape of the entropy curves match favorably with experimental measurements [zhangHeavyFermionsMass2025, rozenEntropicEvidencePomeranchuk2021].
.6 Discussion
The spectrosopic $\pm 10$meV feature, the evolution of entropy with filling and temperature and the pronounced particle–hole asymmetry observed, for example, in charge compressibility measurements constitute key experimental constraints for a proper theoretical description of twisted bilayer graphene. Existing theoretical efforts have partially addressed this set of phenomena using a broad range of frameworks, including low-energy Dirac semimetals [pierceUnconventionalSequenceCorrelated2021, rozenEntropicEvidencePomeranchuk2021], interaction-induced band renormalizations [rademakerChargeSmootheningBand2019, zhuWeakCouplingTheory2024], lifting of flavor degeneracies via spontaneous symmetry breaking or external perturbations [xieWeakFieldHall2021, bultinckGroundStateHidden2020], heavy-fermion and mixed-valence scenarios [zhangHeavyFermionsMass2025, lauTopologicalMixedValence2025], incipient subband formation in paradigmatic Hubbard models [wongCascadeElectronicTransitions2020, ghoshThermopowerProbesEmergent2025], and exotic trion excitations [zhaoTopologicalMottLocalization2025, ledwith2025nonlocalmomentschernbands]. However, a single microscopic framework capable of consistently accounting for all these effects has remained elusive.
Here, we showed that these seemingly distinct experimental signatures can only be understood by simultaneously accounting for correlations, strain and lattice-relaxation effects inherent to realistic twisted bilayer graphene. We identified strain-induced splitting of the flat-band manifold as the central mechanism underlying both the emergence of a persistent feature in the local spectral density and the crossover from an eightfold-degenerate to a fourfold-degenerate local moment upon lowering the temperature [zhangHeavyFermionsMass2025]. At low temperatures, one of the two strain-split sectors becomes effectively inactive and is energetically positioned so as to account quantitatively for the persistent ±10 meV feature observed in QTM and STM experiments. The freezing of this inactive sector naturally explains the reduction of the local-moment degeneracy, and our numerical estimates of the resulting entropy are in excellent agreement with recent experimental measurements.
The identification of an inactive charge sector also has important implications for future theoretical modeling of correlated states in twisted bilayer graphene [songKondoPhaseTwisted2024]. Any faithful low-energy theory must incorporate the effective freezing of one sector upon lowering the temperature and collective instabilities—most prominently superconductivity—must be understood as arising within this constrained low-energy landscape. Finally, lattice relaxation and the associated breaking of particle–hole symmetry render the bonding flat bands more dispersive than the antibonding flat bands; this asymmetry is inherited by the interacting system and naturally leads to stronger correlation effects on the electron-doped side compared to the hole-doped side. More broadly, this work underscores the exceptional sensitivity of correlated quantum materials to externally imposed symmetry breaking, revealing a general route toward controlling collective electronic behavior.
.7 Data availability
The data that support the findings of this work will be deposited in Zenodo and made publicly available upon publication.
.8 Acknowledgements
@@
We thank Oskar Vafek, Erez Berg and Shahal Ilani for useful discussions. L.C., G.R., and T.W. acknowledge support from the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ – EXC 2056 (Project No. 390715994), and SPP 2244 (WE 5342/5-1 project No. 422707584). L.C. gratefully acknowledges the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b158cb. G.R., L.C., R.V., G.S., and T.W. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through QUAST FOR 5249 (Project No. 449872909, projects P4 and P5). G.S. acknowledges financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC2147 “ct.qmat” (project‐id 390858490). R.V. thanks the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the TRR 288 - 422213477 and project Nr. VA 117/23-1 — 509751747. B.A.B. and H. H. were supported by the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070), the Office of Naval Research (ONR Grant No. N00014-20-1-2303), the Global Collaborative Network Grant at Princeton University, the Simons Investigator Grant No. 404513, the NSF-MERSEC (Grant No. MERSEC DMR 2011750), Simons Collaboration on New Frontiers in Superconductivity (SFI-MPS- NFS-00006741-01), the Schmidt Foundation at Princeton University and the Princeton Catalyst Initiative. H. H. and D. C. also acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833).
@@
Supplemental material for Interplay between many-body correlations, strain and lattice relaxation in twisted bilayer graphene
@@
The noninteracting Hamiltonian of the THF model, as originally derived in [songMagicAngleTwistedBilayer2022], has the form
@@
To this Hamiltonian we add the following terms, a subset of those derived in [herzog-arbeitmanHeavyFermionsEfficient2024], which account for uniaxial heterostrain
@@
The THF Hamiltonian involves a sum over moiré lattice sites and and momenta alike. We consider a single lattice site labeled by $\mathbf{R}=(0,0)$, and restrict our $k$-summation on the first mBZ, sampled with a regular centered grid as in [raiDynamicalCorrelationsOrder2024]. The $k$-points on the edges and corners of the hexagonal BZ are degenerate, and are sampled with a relative weight of $1/2$ and $1/3$ respectively. We then strain the obtained $k$-point mesh with the transformation
@@
These are density-density interactions acting within the $f$ subspace, between $f$ and $c$ and within $c$ respectively. The remaining terms, such as the fcf-c exchange interaction, have been neglected since they do not qualitatively alter the resulting picture in the symmetric phase [raiDynamicalCorrelationsOrder2024]. We define the density operators
@@
We assume for simplicity the same interaction coefficient for the $\Gamma_{3}$ and $\Gamma_{1}\oplus\Gamma_{2}$ dispersive orbitals, so that $W_{a}=W=47.12meV$, and $V=48.33meV$. These assumptions are consistent with the study of the unstrained and unrelaxed model in the symmetric phase in [raiDynamicalCorrelationsOrder2024, huSymmetricKondoLattice2023], and do not alter the qualitative picture.
@@
We now show that a similar conclusion can be alternatively reached using symmetry arguments in combination with a perturbative expansion in the strength of the strain. The most general strain term in the THF model contains the isotropic strain parameter $\epsilon_{+}$, the anisotropic $\epsilon_{-}$ and the shear $\epsilon_{xy}$. These transform as $x^{2}+y^{2}$, $x^{2}-y^{2}$ and $xy$ respectively under spatial symmetry. $f$ electron forms the $\Gamma_{3}$ irreducible representation [songMagicAngleTwistedBilayer2022, herzog-arbeitmanHeavyFermionsEfficient2024]. We expand the local hybridization function in powers of the strength of strain, and discuss the symmetry properties of each term. Without loss of generality, we focus on valley $\eta=+$ and spin $s$. The local hybridization can be written as
@@
For the numerical solution of the THF model we employ the w2dynamics software suite [wallerbergerW2dynamicsLocalOne2019], which belongs to the class of DMFT-CTQMC solvers. While extremely robust and able to work with continuous baths and in an extended temperature range, this class of solvers famously suffers from a limitation called fermionic sign problem (see, e.g. Ref. [gullRMP]). No exact prescription for the improvement of the quality of the Monte Carlo sign exists, but it is generally true that off-diagonal local impurity Hamiltonians and hybridization functions tend to aggravate the problem [shinaokaPRB]. In addition to a generally better Monte Carlo sign, another advantage of working with orbital/valley/spin-diagonal impurity problems is the necessity of evaluating a smaller number of matrix elements for the local Green’s function and self-energy. Furthermore, upon postprocessing there is no need to analytically continue the off-diagonal elements of $G$ and $\Sigma$, which is a particularly delicate process.
@@
In the rotated basis, the $f$-subspace consists of doubly-degenerate levels, whose local splitting is $2M_{f}\epsilon_{-}$. This splitting is constant in temperature and linear in strain, as described in [herzog-arbeitmanHeavyFermionsEfficient2024] and visualized in Fig. S2.
@@
We solve the strained and relaxed THF model within the framework of Dynamical Mean-Field theory [georgesDynamicalMeanfieldTheory1996], making use of the Continuous-Time Quantum Monte Carlo (CTQMC) solver w2dynamics [wallerbergerW2dynamicsLocalOne2019]. As discussed in section S1.3, the intraction Hamiltonian is treated at two different levels of approximation: the $H_{U}$ is dynamically accounted for within the DMFT approximation, while the remaining ones are mean-field decoupled. Concordantly, the electronic self-energy of the lattice problem is obtained self-consistently: first, starting from the density matrix at each loop of the DMFT simulation, the mean-field terms are obtained. These are summed to the non-interacting Hamiltonian. The $f$-orbital peojection of this operator represents the impurity Hamiltonian which is solved via CTQMC. The resulting self-energy is upfolded in the combined $f\oplus c$ space by padding it with zeros, since no dynamical correlation effects are considered for the dispersive bands. Self-consistent update of the density matrix and self-energy are then performed until convergence. Since we treat all interaction terms apart from $H_{U}$ at the Hartree level, we can perform a further simplification [huSymmetricKondoLattice2023] that allows us to describe them as an effective “double-counting” term, i.e. an orbital-selective chemical potential acting on the correlated subspace only:
@@
We run the simulations, consistently with [raiDynamicalCorrelationsOrder2024], on 144 CPU cores, using ${\texttt{Nmeas}}=7.5\cdot 10^{4}$ measurements per core for each Monte Carlo simulation, with a number of steps ${\texttt{Ncorr}}=2000$ per core between successive measurements to avoid autocorrelation effects and ${\texttt{Nwarmup}}=2\cdot 10^{6}$ warmup steps per core before each simulation. We measured in imaginary time domain and used the Legendre polynomial basis, with maximum order $\texttt{NLegMax}=40$.
@@
Analytical continuation is performed starting from the last DMFT run using the MaxEnt method from the ana_cont software suite [kaufmannAnaContPython2023]. Once the self-energy and Green’s function are analytically continued, they can be rotated back in the original orbital basis, where $H^{ff}$ is not diagonal. We used a regular mesh of $1001$ real-axis frequency points in the interval [250:+250][-250:+250] meV for $A(\omega)$ and [300:+300][-300:+300] meV for $\Sigma(\omega)$ and no preblur.
@@
Fig.  S6 shows the momentum-resolved spectral function along the $K_{M}\rightarrow\Gamma_{M}\rightarrow-K_{M}\rightarrow M_{M}$ high-symmetry path. The panels in the first row refer to all integer fillings in the range $\nu\in[-4;+4]$ The doping values have been chosen to be directly comparable to the experimental data in [xiaoInteractingEnergyBands2025]. The simulations are in a rather good agreement with the QTM results: they feature flat bands almost everywhere in the mBZ, apart from near the $\Gamma_{M}$ point where they become dispersive. At the CNP, the flat bands are separated by about 35 meV, in agreement with the experiment, and they touch near the $\Gamma_{M}$ point. At positive and negative integer $\nu$ values, a persistence spectral maximum due to the flat band is visible at negative and positive $\omega$ respectively, the value of which is in the range $\pm[10;20]$ meV. Increasing occupation from $\nu=-4$, the lower flat band first remains rather pinned at zero frequency up to $\nu=-1$, after which it shifts to lower frequency and becomes the negative-bias persistent feature for electron doping. Comparing positive and negative doping of the same size, it is interesting to notice how the electron-doped side features a larger spectral gap at large doping with respect to the hole-doped side: this is most evident for $\nu=+4$, where the system is in a clear band-insulating state. This is to be compared to the $\nu=-4$ case, which features a vanishing indirect gap at the Fermi level, and is a product of the $P$-symmetry breaking relaxation terms, which shift the $f$-bands asymmetrically within the hybridization gap. The $\nu=-4$ case is particularly delicate, since the presence or absence of the gap is directly linked to the model parameters, in particular the effective chemical potentials deriving from lattice relaxation.
Figure S7: Occupation with respect to CNP as a function of the chemical potential, total value (a) and orbital-valley resolved (b-c), for T=11.6 K. The blue curves refer to DMFT simulations in presence of strain and lattice relaxation, while the orange one are the unperturbed results from [raiDynamicalCorrelationsOrder2024].
@@
One of the innovations provided by the Quantum Twisting Microscope is the possibility of accessing the spectral function at different points in the Brillouin Zone. By probing $A(\omega)$ at the $\Gamma_{M}$ point, the authors of [xiaoInteractingEnergyBands2025] observe a spectral maximum inside the flat band gap, which evolves with doping in a manner reminiscent of the $\mu(\nu)$ curve (see Fig.1e) in the main text.
@@
In Fig. S9, in the left panel we show the spectral function at the $\Gamma_{M}$ point as a function of doping. This features four wiggling spectral maxima, that evolve from positive to negative frequency as a function of doping. The other panels of Fig. S9 show the momentum-resolved spectral functions for $c$-electrons at five different values of doping, including the CNP. The positions and relative distances of the peaks, are again represented by colored circles and dashed lines. It is immediate to notice that the circles all refer to local spectral maxima of $c$ character. By comparing the relative distances, it is also easy to notice that the four peaks are almost rigidly shifted upon varying the chemical potential. This fact can be understood in the following way: since the $c$ electrons are essentially uncorrelated, their behavior of their spectral distribution is well approximated by that of noninteracting electronic bands. Varying the chemical potential by $\Delta\mu$ has the only effect of rigidly shifting the whole band structure by $\Delta\mu$ with respect to zero frequency. This will follow the peculiar wiggling behavior that results from the $f/c$ spectral weight reshuffling, sometimes referred to as Dirac revivals, as described in [raiDynamicalCorrelationsOrder2024].